# BeFA: A General Behavior-driven Feature Adapter for Multimedia Recommendation

### v1

@@
Abstract.
Multimedia recommender systems focus on utilizing behavioral information and content information to model user preferences. Typically, it employs pre-trained feature encoders to extract content features, then fuses them with behavioral features. However, pre-trained feature encoders often extract features from the entire content simultaneously, including excessive preference-irrelevant details. We speculate that it may result in the extracted features not containing sufficient features to accurately reflect user preferences.
@@ Abstract.
To verify our hypothesis, we introduce an attribution analysis method for visually and intuitively analyzing the content features. The results indicate that certain products’ content features exhibit the issues of information drift and information omission, reducing the expressive ability of features. Building upon this finding, we propose an effective and efficient general Behavior-driven Feature Adapter (BeFA) to tackle these issues. This adapter reconstructs the content feature with the guidance of behavioral information, enabling content features accurately reflecting user preferences. Extensive experiments demonstrate the effectiveness of the adapter across all multimedia recommendation methods. The code will be publicly available upon the paper’s acceptance.
Multimodal Recommendation, Feature Adapter, Information Drift, Information Omission
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06submissionid: 123-A56-BU3
Figure 1. Illustration of content features that do not accurately reflect the users’ preferences. Excessive irrelevant information hinders the recommender system’s ability to effectively model the users’ true preferences.
1. Introduction
@@
Recommender systems have gained widespread adoption across various domains, aiming to assist users in discovering information that aligns with their preferences (vldb1, ; vldb2, ). In the case of multimedia platforms, the abundance of data resources provides recommender systems with increased opportunities to accurately model user preferences (survey, ).
@@
Existing multimedia recommendation methods typically involves two main steps (w2go, ; survey, ). First, a pre-trained feature encoder is employed to capture content features from diverse modalities. Subsequently, these content features are fused with behavioral features to obtain user preference representations. Recently, researchers have focused on enhancing the quality of these representations through self-supervised learning. For instance, SLMRec (slmrec, ) investigates potential relationships between modalities through the use of contrastive learning, thereby obtaining a powerful representation. BM3 (bm3, ) utilizes a dropout strategy to construct multiple views and reconstructs the interaction graph, incorporating intra- and inter-modality contrastive loss to facilitate effective representation learning. MICRO (micro, ) and MGCN (mgcn, ) maximize the mutual information between content features and behavioral features with a self-supervised auxiliary task, and have achieved excellent performance.
Figure 2. Results of visualisation attribution analysis on the TMALL dataset.(a) is the original product image and (b) is its corresponding heatmap.
@@
Despite the good progress made by existing methods in utilizing content and behavioral information more effectively, a crucial yet easily overlooked problem arises: Are content features obtained from pre-trained encoders containing sufficient features to reflect user preferences? Intuitively, multimedia content inherently exhibits the characteristic of low informational value density, where a significant portion of the presented information may be irrelevant to the users’ focus (micro, ; survey, ). Pre-trained feature encoders extract information from the entire content simultaneously, which can result in content features that do not truly reflect the users’ preferences (as shown in Figure 1). Fusing these irrelevant content features with behavioral features may mislead user preference modeling, resulting in suboptimal recommendation performance.
@@
To answer this question, we introduce a similarity-based attribution analysis method for visualizing and intuitively analyzing the content features. This method evaluates the extent to which content features can reflect user preferences, enabling researchers to visually assess the quality of content features for the first time. The results indicate that not all products’ content features accurately reflect user preferences. Due to the presence of irrelevant information, certain products’ content features exhibit the issues of information drift and information omission. As shown in Figure 2, some products’ content features do not include information about the products that users are interested in, but instead erroneously include information about unrelated items, a phenomenon we term information drift. There are also some products’ content features that omit certain key details of the products, which we refer to as information omission. These issues ultimately prevent recommender systems from accurately modeling user preferences. To verify the rationality of these findings, we provide a theoretical analysis and explanation. Furthermore, we propose a plug-and-play general Behavior-driven Feature Adapter (BeFA) to address the discovered issues. This adapter effectively decouples, filters and reconstructs content features, leveraging behavioral information as a guide to obtain more precise representations of content information. Extensive experiments demonstrate the adapter’s effectiveness across various multimedia recommendation methods and feature encoders.
@@
  • We experimentally revealed the issues of information drift and information omission in content features. We also provide a theoretical analysis to validate the rationality of these findings.
@@
Figure 3. Pipeline of the proposed attribution analysis method.
2. RELATED works
2.1. Multimedia Recommendation
Collaborative filtering (CF)-based approaches have achieved great success in recommender systems, relying on behavioral similarities to make top-k recommendations (cf1, ; cf2, ). Since user preferences are usually influenced by multimodal information, it facilitates researchers to integrate multimodal content information into CF-based approaches with the aim of improvement. Typically, content features are first extracted using pre-trained neural networks and then fused with behavioral features to enhance preference modeling (vbpr, ; w2go, ). Considering that behavioral information can be naturally modeled as a bipartite graph structure, researchers begin work on recommender systems based on Graph Convolutional Network(GCN) (gcn1, ; gcn2, ). For instance, MMGCN citemmgcn constructs modality-specific user-item interaction graphs to model user preferences specific to each modality. LATTICE (lattice, ) adds links between items with similar modality characteristics and create a separate item-item graph for each modality. FREEDOM (FREEDOM, ) constructs item-item graphs to help learn item semantic relations and reduces the graphs according to the sensitivity edge pruning technique. Meanwhile, researchers began to focus on enhancing the quality of representations through self-supervised learning. BM3 (bm3, ) learns user and item representations by reconstructing the user-item interaction graph and aligning modality features from both inter and intra-modality perspectives. MGCN (mgcn, ) comprehensively models user preferences by adaptively learning the relative importance of different modality features.
However, existing methods usually directly utilize features extracted from pre-trained feature encoders as side information for each item. The pre-trained feature encoders are trained on large-scale datasets collected from the natural world. The features of natural data and e-commerce data are very different. In e-commerce scenarios, the content information often contains a lot of irrelevant information and exhibits the characteristic of low value density. For example, only a small portion of the content image contain relevant information, while the rest is considered noise. This disparity leads to a substantial amount of noise in the contetn features extracted by pre-trained encoders when applied to multimodal recommender systems. Consequently, the content features do not contain enough features to sufficiently reflect users’ preferences. Fusing these irrelevant content features with behavioral features may mislead user preference modeling, resulting in suboptimal recommendation performance.
2.2. Attribution Analysis
Methods for modality feature attribution analysis have been widely studied, aiding in the visualization of modality feature content. Class Activation Mapping (CAM) (CAM, ) is a method used to visualize the decision-making process of convolutional neural networks by generating heatmaps that highlight regions of an image contributing to a particular class prediction. Grad-CAM (gradcam, ) enhances this by using the gradient of the target class with respect to the feature maps of a specific convolutional layer, creating heatmaps that can be applied to various network architectures without altering the network structure. Score-CAM (scoreCAM, ) further improves upon Grad-CAM by weighting the activation map using output scores instead of gradient information, thus avoiding gradient-related problems. In multimodal recommender systems, t-SNE (tsne, ) is often used to visualize the distribution of samples in high-dimensional feature spaces, providing insights into the relationships and distributions among different modality features.
However, current methods (cvpr, ; gScoreCAM, ) for analyzing image feature weights predominantly rely on CAM-based methods. The heatmaps generated by these methods only reflect the contribution of certain image regions to classification tasks or the relevance of text features, failing to capture the role of multimodal features in recommender systems. Most existing multimodal recommendation works use use t-SNE (tsne, ) to analyze the modality features because t-SNE can visualize multimodal feature distributions. But t-SNE can only show the distribution of features in 2-dimensional space, which lacks the intuitive effectiveness of CAM-generated heatmaps and is not suitable for understanding and analyzing specific content features. Consequently, existing attribution analysis methods exhibit significant deficiencies in analyzing multimodal features within multimodal recommender systems.
2.3. Parameter-Efficient Adaptation
In the field of computer vision (CV) and natural language processing (NLP), several studies have identified that semantic differences can restrict the expressiveness of features, thereby impacting the model’s performance as well as downstream tasks (nlp1, ; cv1, ). To address this challenge, researchers have explored techniques such as fine-tuning and adapters to bridge these semantic differences (nlp2, ; softprompt, ; lora, ). However, in the context of recommendation scenarios where data is continuously changing and expanding, performing full fine-tuning of the encoder poses significant challenges. It often requires a significant amount of time and computational resources, and it may even lead to a decline in the model’s ability to generalize to new data. Given these practical limitations, retraining the feature encoder from scratch for every change in the data is unrealistic. Most of the existing multimodal recommendation methods directly use pre-trained features without fine-tuning, but the content feature extracted by the pre-trained feature encoders may not be sufficient to reflect the users’ preferences, so we consider introducing feature adapters. One such approach is Low-Rank Adaptation(LoRA) (lora, ), which reduces the number of trainable parameters by integrating trainable low-rank decomposition matrices into the Transformer architecture. Another approach is Prompt Tuning (softprompt, ), which offers a different perspective by incorporating learnable embedding vectors, acting as hints, into the inputs of the pre-trained model.
However, adding adapters to the middle layer of the pre-trained model may lead to some information loss. Existing parameter tuning methods are not generalized for recommendation tasks, as they fail to adequately incorporate behavioral information, which intuitively reflects the users’ preferences and is crucial for an effective recommender systems.
3. PRELIMINARIES
In this section, we analyze and prove in detail the problems of the content features extracted by pre-trained encoders for recommendation tasks. In section 3.1 we introduce an attribution analysis for visual attribution analysis of content features in multimodal recommender systems. In section 3.2, we specifically analyze the deficiencies of pre-trained feature encoders. In section 3.3 we analyze and prove the problems from a mathematical and theoretical point of view.
3.1. Attribution Analysis
Existing methods for analyzing the weights of image features are predominantly based on class activation mapping (CAM) (CAM, ). However, the heatmaps generated by CAM only reflect the contribution of certain image regions to the classification task and fail to assess their impact on recommendation models (scoreCAM, ; gScoreCAM, ). These limitations restrict our in-depth understanding of modality feature processing in recommender systems. Thus we designed the attribution analysis method for analyzing image features in multimodal recommendations based on the characteristics of recommender systems.
To visually illustrate the issues raised, we introduce a intuitively analysis for the attribution of multimodal features in the recommender system. This method aims to explore the effectiveness of content features extracted by pre-trained multimodal feature encoders in multimodal recommender systems. It also provides a visual demonstration of the effectiveness of our proposed adapter in bridging the gap between the content features extracted by the pre-trained encoders and the features required by the recommender system. To accurately identify which parts of the content features users pay attention to, we use cosine similarity as a measurement. By calculating the cosine similarity of each pixel to the behavioral features in the recommendation models, we generate a saliency heatmap. This heatmap is created by weighting and summing the masked feature maps based on the calculated similarities. This approach visualizes the portions of the content features that are truly effective in the recommendation task, providing a deeper understanding of the role of multimodal features in recommender systems.
(1) $\text{similarity}(x,y)=\frac{{v_{x}\cdot v_{y}}}{{\|v_{x}\|\cdot\|v_{y}\|}}$
Similar to the concept behind CAM-based approaches, we choose the target layer for visualization to ensure interpretability (scoreCAM, ; gScoreCAM, ). In the case of image encoders, for both ResNet (resnet, ) and ViT (ViT, ) versions, we select the channels closest to the CLIP prediction layer. In convolutional neural networks (CNNs), as the layers go deeper, the information represented in the feature maps becomes more abstract and high-level. The feature maps of the last ReLU layer capture the most important high-level features in the input image. ReLU-activated feature maps contain only positive or zero activations, which helps to emphasize the regions that the model considers more relevant. For the ViT series, as the gradients are zero for all layers except for the [CLS] token and only the [CLS] token is used for final predictions in the last layer embedding, we choose the penultimate layer. Consequently for the ResNet series, we select the last ReLU layer of the final Bottleneck as the target layer, for the ViT series, we use the second-to-last ResidualAttentionBlock as the target layer.
In the first stage, we perform a forward pass of the product images through CLIP, obtaining $N$ channels at the target layer denoted as ${A_{i}}^{N}$. The purpose is to discriminate the contributions of different targets across different channels. Subsequently, for each channel, we upsample it, denoted as $A_{i}^{up}$. This is done to generate masked images corresponding to different parts and to calculate the similarity weights of different regions. In the second stage, element-wise multiplication is performed between each mask $A_{i}^{up}$ and all color channels $x$ of the input image, and then the result is compared with the corresponding to behavioral features to compute the similarity. We name our method Behavior Information CAM (BeCAM) and its saliency map is computed by:
(2) $M_{\mathrm{BeCAM}}^{i}=\sum_{i}^{N}A_{i}\times\text{similarity}((x\odot A_{i}^{up}),{ID}_{i})$
where $\odot$ is the Hadamard product.
3.2. Deficiency Analysis
Figure 4. An illustration of a theoretical analysis of deficiency analysis
@@ 3.2. Deficiency Analysis
Figure 5. Comparison of efficient parameter tuning methods. (a) Low-Rank Adaptation (b) Soft Prompt Turning (c) Our proposed adapter.
By analyzing the generated heatmap (shown in Figure 2), it is clear that the pre-trained feature encoder suffers from the issues of wrong region of interest and insufficient region of interest on the product dataset. This can be visualized by the fact that the content features extracted by the pre-trained feature encoders have the problems of information drift and information omission. This is a very visible demonstration of the low information value density of the multimedia content itself, resulting the content features are insufficient to reflect users’ preferences.
@@
Specifically, as shown in Figure 2, although the user wants to buy clothes, the feature encoder incorrectly focuses on the model’s face, resulting in information drift. There are also products whose content features miss some key details of the product, such as only focusing on certain details of the clothes and failing to focus on its whole, resulting in information omission. Information drift can cause the recommender system to ignore useful and critical information. This leads to a loss of important guidance for the recommendation task. It also results in a mismatch between recommendations and users’ preferences. Information omission can result in important visual or textual cues being ignored or incomplete. This causes a loss of critical information needed for the recommendation task. These issues collectively compromise the overall effectiveness of the recommender system. As a result, directly applying the content features extracted by the pre-trained encoder in the multimodal recommender system will inevitably impair the performance of the recommender system.
3.3. Theoretical Analysis
In this section, we reveal the deficiencies of content features extracted by pre-trained feature encoders within the embedding space. We analyze the gap between these content features and the ideal features required by the recommendation model. Furthermore, we demonstrate that optimizing the consistency between the extracted features and the ideal features necessary for accurate recommendation predictions can significantly enhance recommendation performance. This analysis provides theoretical support for the effectiveness of our proposed adapter.
The embedding space captures the semantic relationships and similarities between features. Analyzing features in the embedding space enables a better understanding and interpretation of their relationships. Formally, considering a node ${i}$ and its representation $\boldsymbol{e_{i}}$ in the embedding space, which represents the ideal features required by the recommendation model. On the basis of the item feature representation $\boldsymbol{e_{i}}$ in the embedding space, the recommender system calculates the probability distribution of the rating or clicking behavior of the user $u$ on the item $i$, which can be viewed as the prediction $P(r_{ui}|\boldsymbol{e_{u}},\boldsymbol{e_{i}})$. Where $r_{ui}$ stands for the rating or click behavior of the user $u$ on item $i$, $\boldsymbol{e_{u}}$ represents the user’s feature, and $\boldsymbol{e_{i}}$ signifies the item’s feature. Our objective is to find the optimal user feature $e_{u}$ and item feature $\boldsymbol{e_{i}}$ to maximize the posterior distribution $P(r_{ui}|\boldsymbol{e_{u}},\boldsymbol{e_{i}})$. In other words, we aim to maximize the posterior distribution for the user feature $\boldsymbol{e_{u}}$ and item feature $\boldsymbol{e_{i}}$ parameters. This can be formally described by the following objective function:
(3) $L(\boldsymbol{e_{u}},\boldsymbol{e_{i}})=\arg\max_{\boldsymbol{x_{u}},\boldsymbol{e_{i}}}P(r_{ui}|\boldsymbol{e_{u}},\boldsymbol{e_{i}})$
The representation $\boldsymbol{e_{i}^{\prime}}$ extracted by the feature encoder can be seen as the prior distribution. However, due to semantic differences between the natural data and e-commerce data, there is inevitably some deviations between the features extracted by the pre-trained feature encoder and the ideal features. Let us denote this deviation as $\theta$.
(4) $\theta=\arccos\left(\frac{\boldsymbol{e_{i}}\cdot\boldsymbol{e_{i}^{\prime}}}{\|\boldsymbol{e_{i}}\|\|\boldsymbol{e_{i}^{\prime}}\|}\right)$
Analyzing $\theta$ helps understand the deviation between the features extracted by the feature encoder and the ideal features. A smaller $\theta$ indicates that the two representations are very close in the embedding space. This implies a high consistency between the features extracted by the feature encoder and the ideal features, which benefits the recommendation system in accurately capturing user interests and item similarities. By reducing $\theta$, the features become closer to the ideal features, thereby enhancing the performance of the recommendation system.
The part $p$ of the representation $\boldsymbol{e_{i}^{\prime}}$ extracted by the encoder that is truly effective for the recommendation task can be represented as follows:
(5) $p=\left(\boldsymbol{e_{i}^{\prime}}\cdot\frac{\boldsymbol{e_{i}}}{\|\boldsymbol{e_{i}}\|}\right)\frac{\boldsymbol{e_{i}}}{\|\boldsymbol{e_{i}}\|}=\boldsymbol{e_{i}^{\prime}}\cdot\cos(\theta)$
When the deviation $\theta$ is large, as $\boldsymbol{e_{i}^{\prime}}$ in Figure 4, content features exhibit the issues of information omission, causing the effective length of $p$ deviates significantly from $\boldsymbol{e_{i}}$. This results in a lack of crucial information in the extracted features, reflected as insufficient region of interest in the heatmap. Conversely, when $\theta$ is too large,as $\boldsymbol{e_{j}}$ in Figure 4, content features exhibit the issues of information drift, with $p$ being located in the wrong quadrant and its effective direction opposing $\boldsymbol{e_{i}}$. This introduces a large amount of incorrect information, reflected as incorrect region of interest in the heatmap.
In such circumstances, the recommendation system is easily to make incorrect recommendations, because the feature representation contradicts the users’ preferences, failing to appropriately reflect the characteristics of items or users. In the recommendation task, we aim for a smaller deviation between $\boldsymbol{e_{i}^{\prime}}$ and $\boldsymbol{e_{i}}$. To formalize this objective, we define an error function $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})$ to measure the gap between the representations:
(6) $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})=\left(1-\frac{{\boldsymbol{e_{i}^{\prime}}\cdot\boldsymbol{e_{i}}}}{{\|\boldsymbol{e_{i}^{\prime}}\|\|\boldsymbol{e_{i}}\|}}\right)$
(7) $\Delta=\mathbb{E}_{P(\boldsymbol{e_{i}^{\prime}})}[f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})]$
The function $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})$ measures the deviation between representations. In recommendation tasks, the quality of the feature representation $\boldsymbol{e_{i}^{\prime}}$ directly affects the accuracy and effectiveness of recommendations. A decline in the quality of $\boldsymbol{e_{i}^{\prime}}$, indicated by an increase in the value of the error function $f(\boldsymbol{e_{i}^{\prime}},\boldsymbol{e_{i}})$, donates an augmented diversity between the representation extracted by the pre-trained encoder $\boldsymbol{e_{i}^{\prime}}$ and the ideal representation $\boldsymbol{e_{i}}$. Consequently, this leads to an increase in the expected deviation $\Delta$, subsequently influencing the posterior distribution $P(r_{ui}|\boldsymbol{e_{u},e_{i}})$. Such circumstances may impair the recommendation system’s ability to accurately capturing the users’ preferences, thereby reducing the accuracy and efficacy of recommendations. By minimizing $\Delta$, we can ensure that the representation $\boldsymbol{e_{i}^{\prime}}$ are closer to the ideal representation $\boldsymbol{e_{i}}$ in expectation. This adaptation of content features enhances the consistency with the ideal representation and thus improves the performance of the recommender system.
Figure 6. Visual analysis of the effect of adapter on feature purification. (a) is the original image. (b) is the similarity heatmap of the visual features extracted by the pre-trained CLIP. (c) (d) (e) are the similarity heatmaps of the visual features adapted by BeFA applied on FREEDOM, BM3 and LATTICE respectively
Table 1. Statistics of the experimental datasets
Dataset #User #Item #Behavior Density
TMALL 13,104 7,848 151,928 0.148%
Microlens 46,420 14,079 332,730 0.051%
H&M 43,543 16,915 369,945 0.050%
4. feature adapter
@@ 3.3. Theoretical Analysis
To address the discovered issues, we proposed Behavior-driven Feature Adapter (BeFA). This is a adapter for adapting multimodal content features that are employed by multimodal recommender systems. Specifically, the content features are first extracted by a pre-trained encoder, and then the content features are adapted by BeFA, and then fed into the multimodal recommender system for item and user modeling. BeFA and the downstream recommender system share the optimization objective and use End-to-End training.
4.1. Problem Formulation
Let $u\in\mathcal{U}$ and $i\in\mathcal{I}$ denote the user and item, respectively. The input behavioral features for user $u$ and item $i$ are represented as $\mathrm{E_{id}}\in\mathbb{R}^{d\times(|U|+|I|)}$, where $d$ is the embedding dimension. Each item modality feature is denoted as $\mathrm{E_{i,m}}\in\mathbb{R}^{d_{m}\times|I|}$, where $d_{m}$ is the dimension of the features, $m\in\mathcal{M}$ represents the modality, and $\mathcal{M}$ is the set of modalities. In this paper, we primarily focus on visual and textual modalities, denoted by =$\mathcal{M}=\{v,t\}$. It’s worth noting that our approach is adaptable to incorporate more than two modalities.
Next, the users’ historical behavior feature is represented as $R\in\{0,1\}^{|U|\times|I|}$, where each entry $R_{u,i}=1$ if user $u$ clicked item $i$, otherwise $R_{u,i}=0$ This historical interaction data $R$ can naturally be interpreted as a user-item bipartite graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. In $\mathcal{G}$ , the vertex set $\mathcal{V}=\mathcal{U}\cup\mathcal{I}$ donates all items and users. The edge set =$\mathcal{E}=\{(u,i)|u\in\mathcal{U},i\in\mathcal{I},R_{u,i}=1\}$ represents the interacted users and items. The objective of multimedia recommendation is to accurately predict users’ preferences by ranking items for each user based on predicted preference scores $\hat{y}_{u,i}$.
Table 2. Performance Comparison on Different Recommender Models. The t-tests validate the significance of performance improvements with p-value $\leq$ 0.05.

### v2

@@
Abstract
@@
Multimedia recommender systems focus on utilizing behavioral information and content information to model user preferences. Typically, it employs pre-trained feature encoders to extract content features, then fuses them with behavioral features. However, pre-trained feature encoders often extract features from the entire content simultaneously, including excessive preference-irrelevant details. We speculate that it may result in the extracted features not containing sufficient features to accurately reflect user preferences. To verify our hypothesis, we introduce an attribution analysis method for visually and intuitively analyzing the content features. The results indicate that certain items’ content features exhibit the issues of information drift and information omission, reducing the expressive ability of features. Building upon this finding, we propose an effective and efficient general Behavior-driven Feature Adapter (BeFA) to tackle these issues. This adapter reconstructs the content feature with the guidance of behavioral information, enabling content features accurately reflecting user preferences. Extensive experiments demonstrate the effectiveness of the adapter across all multimedia recommendation methods. Our code is made publicly available on https://github.com/fqldom/BeFA.
Introduction
@@
Recommender systems have gained widespread adoption across various domains, aiming to assist users in discovering information that aligns with their preferences (Parchas et al. 2020; Liu et al. 2017). In the case of multimedia platforms, the abundance of data resources provides recommender systems with increased opportunities to accurately model user preferences (Zhou et al. 2023a).
@@
Existing multimedia recommendation methods typically involves two main steps (Yuan et al. 2023; Zhou et al. 2023a). First, a pre-trained feature encoder is employed to capture content features from diverse modalities. Subsequently, these content features are fused with behavioral features to obtain user preference representations. Recently, researchers have focused on enhancing the quality of these representations through self-supervised learning. For instance, SLMRec (Tao et al. 2023) investigates potential relationships between modalities through the use of contrastive learning, thereby obtaining a powerful representation. BM3 (Zhou et al. 2023c) utilizes a dropout strategy to construct multiple views and reconstructs the interaction graph, incorporating intra- and inter-modality contrastive loss to facilitate effective representation learning. MICRO (Zhang et al. 2022) and MGCN (Yu et al. 2023) maximize the mutual information between content features and behavioral features with a self-supervised auxiliary task, and have achieved excellent performance.
Figure 1: Illustration of content features that do not accurately reflect the users’ preferences. Excessive irrelevant information hinders the recommender system’s ability to effectively model the users’ true preferences.
Figure 2: Results of visualisation attribution analysis on the TMALL dataset. The first row contains the original item images while the second row displays the corresponding heatmaps. The four samples on the left reflect information drift and the four samples on the right reflect information omission.
@@
Despite the good progress made in utilizing content and behavioral information more effectively, a crucial yet easily overlooked problem arises: Are content features obtained from pre-trained encoders containing sufficient features to reflect user preferences? Intuitively, multimedia content inherently exhibits the characteristic of low informational value density, where a significant portion of the presented information may be irrelevant to the users’ focus (Zhou et al. 2023a). Pre-trained feature encoders extract information from the entire content simultaneously, which can result in content features that do not truly reflect the users’ preferences (as shown in Figure 1). Fusing these irrelevant content features with behavioral features may mislead user preference modeling, resulting in suboptimal recommendation performance.
@@
To answer this question, we introduce a similarity-based attribution analysis method for visualizing and intuitively analyzing the content features. This method evaluates the extent to which content features can reflect user preferences, enabling researchers to visually assess the quality of content features for the first time. The results indicate that not all items’ content features accurately reflect user preferences. Due to the presence of irrelevant information, certain items’ content features exhibit the issues of information drift and information omission. As shown in Figure 2, some items’ content features do not include information about the items that users are interested in, but instead erroneously include information about unrelated items, a phenomenon we term information drift. There are also some items’ content features that omit certain key details of the items, which we refer to as information omission. These issues ultimately prevent recommender systems from accurately modeling user preferences. Furthermore, we propose a plug-and-play general Behavior-driven Feature Adapter (BeFA) to address the discovered issues. This adapter effectively decouples, filters and reconstructs content features, leveraging behavioral information as a guide to obtain more precise representations of content information. Extensive experiments demonstrate the adapter’s effectiveness across various recommendation methods and feature encoders.
@@
  • We experimentally revealed the issues of information drift and information omission in content features.
@@
Related Work
Figure 3: Pipeline of the proposed attribution analysis method.
Multimedia Recommendation
Collaborative filtering (CF) based methods leverage behavioral similarities for top-k recommendations (Zhang et al. 2014). To improve the performance of CF-based methods, researchers integrate item multimodal content. Typically, they employ pre-trained neural networks to extract content features, and then fuse content features with behavioral features to more effectively model user preference. (Yuan et al. 2023). For instance, MMGCN (Wei et al. 2019) creates modality-specific interaction graphs. LATTICE (Zhang et al. 2021) adds links between similar items. FREEDOM (Zhou and Shen 2023) constructs graphs using sensitivity edge pruning. Recently, self-supervised learning methods like BM3 (Zhou et al. 2023c) and MGCN (Yu et al. 2023) further enhance representation quality by reconstructing interaction graphs and adaptively learning content feature importance. However, item content information often contains much irrelevant information and exhibiting low value density. We argue pre-extracted item content features insufficiently reflect users’ preferences. Directly incorporating these features may mislead user preference modeling, resulting in suboptimal recommendation performance.
Attribution Analysis
Modality feature attribution analysis methods such as Class Activation Mapping (CAM) (Zhou et al. 2016), Grad-CAM (Selvaraju et al. 2017), and Score-CAM (Wang et al. 2020) generate heatmaps to visualize the decision-making process of convolutional neural networks. Grad-CAM uses gradients for better applicability across network architectures. Score-CAM weight activation maps with output scores, which eliminates the dependence on gradients. However, current attribution analysis methods primarily focus on analyzing the encoder itself and fail to visually analyze items’ content features in recommender systems. Most of these methods use gradient information to obtain weights, which reflect more of the encoder focus. As a result, they fail to capture the similarity between content features and behavioral features, which directly reflect the user’s preferences. Consequently, these methods are unsuitable for use in recommender systems. Thus, to better analyze the quality of item content features, we need to design a novel similarity-based attribution analysis method.
Parameter-Efficient Adaptation
In multimodal domains, semantic differences limit feature expressiveness, impacting model performance (Khurana et al. 2023; Rinaldi, Russo, and Tommasino 2023). Using pre-trained encoders to extract features in real-world scenarios can result in not accurately reflecting user preferences. To address this challenge, researchers have explored techniques such as fine-tuning and adapters to reduce the impact of semantic differences. In the context of recommendation scenarios where data is continuously changing and expanding, performing full fine-tuning of the encoder faces huge challenges. Due to the above limitations, we consider low-parameter fine-tuning methods to efficiently adapt content features in real time. Low-Rank Adaptation (LoRA) (Hu et al. 2021) integrate low-rank decomposition matrices, while Prompt Tuning (Lester, Al-Rfou, and Constant 2021) uses learnable embedding vectors as hints. However, adding adapters to the middle layer of the pre-trained model may lead to information loss. Existing methods are not generalized for recommendation tasks. They fail to adequately incorporate behavioral information, which is crucial for intuitively reflecting users’ preferences.
Preliminaries
Attribution Analysis
Existing attribution methods are mainly focused on pre-trained encoders, which can only analyze the quality of the content features themselves, fail to reflect their contributions to recommendations. Considering that behavioral features directly reflect user preferences in recommendation, we propose a similarity-based attribution analysis. This method explores the quality of content features extracted by a pre-trained encoder in recommendations. By calculating the cosine similarity of each pixel to the behavioral features, we generate a heatmap that visually identifies the content features relevant to users’ behaviors.
We follow the core principle of Class Activation Mapping (CAM) to linearly weight feature maps. The main difference compared to the previous series of CAM methods is the way in which the linear weights are obtained. Our approach consists of two primary stages. In the first stage, we take the item image $I$ as input to the corresponding encoder, obtaining $N$ images on the corresponding channels at the target layer denoted as ${a_{l}}$. Then we upsample it, denoted as $a_{l}^{up}$. This is done to generate masked images corresponding to different parts. Considering in all types of encoders, layers closer to the prediction layer contain the highest level and most abstract information (Vaswani et al. 2017). Consequently, for the ResNet family visual encoder, we select the last ReLU layer of the final Bottleneck as the target layer, while for the ViT family visual encoder, we choose the penultimate ResidualAttentionBlock. The corresponding content feature $\boldsymbol{x_{l}}$ is extracted from each mask $a_{l}^{up}$. Behavioural features $\boldsymbol{h_{i}}$ are obtained from existing multimedia recommendation models. The weights of the corresponding content features are obtained by calculating the similarity between them.
@@
Figure 4: Comparison of parameter tuning methods. (a) Low-Rank Adaptation (b) Soft Prompt Turning (c) BeFA.
$\text{similarity}(\boldsymbol{x_{l}},\boldsymbol{h_{i}})=\frac{{\boldsymbol{x_{l}}\cdot\boldsymbol{h_{i}}}}{{\|\boldsymbol{x_{l}}\|\cdot\|\boldsymbol{h_{i}}\|}}$ (1)
Ultimately, the final visualization result is achieved by linearly weighting and summing the feature maps with the similarity. The heatmap $M$ is represented as:
$M=\sum_{l}^{N}a_{l}\times\text{similarity}(f_{pre}(I_{c}\odot a_{l}^{up}),\boldsymbol{h_{i}})$ (2)
where $\odot$ is the Hadamard product. $I_{c}$ are the all color channels of the image. $f_{pre}$ denotes the process of extracting features using the corresponding encoder.
Deficiency Analysis
@@
From Figure 2 we observed that the pre-trained feature encoder suffers from the issues of wrong region of interest (information drift) and insufficient region of interest (information omission). Specifically, although the user wants to buy clothes, the feature encoder incorrectly focuses on the face of the person who wears the cloth, a phenomenon we call information drift. Meanwhile, we also found that some of the features did not fully reflect all the characteristics of the item. Such as only focusing on certain details of the clothes and failing to focus on its whole, a phenomenon we call information omission. Information drift can cause the recommender system to ignore useful and critical information. This leads to a loss of important guidance for the recommendation task. Which lead a mismatch between recommendations and users’ preferences. Information omission can result in important visual or textual cues being ignored or incomplete. This causes a loss of critical information needed for the recommendation task. These issues compromise the overall performance of the recommender system. As a result, directly applying the pre-extracted content features will inevitably impair the performance of the recommender system. In the Appendix, we perform a theoretical analysis to further illustrate the point.
Feature Adapter
@@
To address the discovered issues, we propose Behavior-driven Feature Adapter (BeFA). The purpose of this adapter is to efficiently adapt content features to improve their quality in recommendations, thus improving the performance of recommendations. Specifically, the content features are first extracted by a pre-trained encoder and then adapted by BeFA. Then it is fed into the multimodal recommender system for item and user modeling. BeFA and the downstream recommender system share the optimization objective and adopt End-to-End training strategy.
Problem Formulation